OpenStudy (anonymous):

I need help with this math question can someone please help me? If f(x)=2x^2-4x-3 and g(x)=(x-4)(x+4),find f(x)-g(x) Answers: 4x+13 x^2-4x+13 3X^2-4x+13 -4x^2-4x+13

OpenStudy (welshfella):

f(x) - g(x) = 2x^2 - 4x - 3 - (x-4)(x + 4) first expand (x - 4)(x + 4) then subtract

OpenStudy (anonymous):

How do I do that? @welshfella

OpenStudy (welshfella):

(x + 4(x - 4) = x(x -4)+4(x - 4)

OpenStudy (anonymous):

So is it like the foil method?? no? @welshfella

OpenStudy (welshfella):

you can use the foil method if you prefer it

OpenStudy (anonymous):

oh okay, so is the answer -4x^2-4x+13??

OpenStudy (welshfella):

No tell me what you get for (x+4)(x-4)

OpenStudy (anonymous):

okay is it this \[x^2-16+4x+4x\]?

OpenStudy (anonymous):

kind of what i got when i foiled them... as you can see im not good in math at all @welshfella

OpenStudy (welshfella):

nearly write but its +4x - 4x

OpenStudy (anonymous):

okay i see

OpenStudy (welshfella):

so the result is x^2 - 16 so now we have 2x^2 - 4x - 3 - (x^2 - 16) Note the parentheses around the x^2 - 16 Can you simplify this?

OpenStudy (anonymous):

no i dont think so

OpenStudy (welshfella):

you have to distribute the negative over the parentheses so -16 becomes -1*-16 = + 16

OpenStudy (anonymous):

you completely lost me there 0_o

OpenStudy (welshfella):

2x^2 - 4x - 3 - (x^2 - 16) = 2x^2 - 4x - 3 - x^2 + 16 now simplify like terms and you have the answzer

OpenStudy (anonymous):

3x^2-4x+13??

OpenStudy (welshfella):

bring like terms together to make it easier 2x^2 - x^2 - 4x -3 + 16

OpenStudy (welshfella):

what is 2x^2 - x^2 ?

OpenStudy (welshfella):

i know this sounds childish but suppose x^2 = an apple what we have here is 2 apples take away one apple

OpenStudy (welshfella):

leaving one apple, right?

OpenStudy (anonymous):

right

OpenStudy (welshfella):

2 apples - 1 apple = 1 apple 2x^2 - 1 x^2 = ?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

?

OpenStudy (welshfella):

no x^2 = apple instead of an apple write x^2

OpenStudy (anonymous):

wait so is it x^2-4x+13???

OpenStudy (welshfella):

yes 2 x^2 - x^2 = x^2 we could write 1 x^2 but the 1 is usually left out 2x^2 and x^2 are like terms so you can subtract the

OpenStudy (anonymous):

omg thank you so much for all of the help!! and for not giving up on me on this question @welshfella

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

think of algebra as a kind of shorthand will help remember you can only simplify like terms 4x + x^2 cann ot be simplified because they are not like terms but 2x and 5x can - 2x + 5x = 7x

OpenStudy (welshfella):

* 2x + 5x = 7x ( i didnt mean to put the negative there)

OpenStudy (anonymous):

Thank you for that tip i will use it with any other math questions @welshfella

OpenStudy (anonymous):

if you dont mind could you help me with more? @welshfella

OpenStudy (welshfella):

sorry I have to go right now